OpenStudy (anonymous):

Write the formula for the geometric sequence. 5, 10, 20, 40, …

OpenStudy (anonymous):

A. an=40∙2n-1 B. an=40∙2n C. an=5∙2n-1 D. an=5∙2n

OpenStudy (***[isuru]***):

can u find the common ratio ?

OpenStudy (anonymous):

And for the first three. A. B. and C. the an is a^n but the n is on th bottom not the top. does that make scene?

OpenStudy (anonymous):

5(nx2) where n is the place (1st number, 2nd number etc) answer is D

OpenStudy (***[isuru]***):

if we consider the 2 consecutive terms of a geometric progression a, b the common ration (r) would be \[r = \frac{ b }{ a }\]

OpenStudy (***[isuru]***):

Auxuris u r wrong

OpenStudy (***[isuru]***):

and dont just give the answer

OpenStudy (***[isuru]***):

the formula of a geometric progression is \[T_{n} = ar^{n-1}\]

OpenStudy (anonymous):

So @***[ISURU]*** umm.... ok sorry but im really bad at math. and im confused!

OpenStudy (anonymous):

Dont know why that just posted.

OpenStudy (anonymous):

so what do we have to figure out first.

OpenStudy (anonymous):

oh oops tried an example and figured that you'll have to subtract 1 because it starts with a 5 hang on, shouldn't it be 5x(n-1)2?

OpenStudy (anonymous):

wait what? @***[ISURU]*** @Auxuris you guys! I need help please? Anyone?